OpenStudy (anonymous):

Point slope, slope intercept, and standard form of the equation of a line that passes through the points (a1, b1) and (a2, b2)?

OpenStudy (anonymous):

find the gradient m= b1-b2/a1-a2 then the y -y1 =m (x - x1) in your case its y - b1 = m(x-a1) y = m(x-a1)+b1